The loan of Php1,350 is to be repaid in 1 year at 4.5% simple interest. How much is added to the loan?

1
Expert's answer
2021-11-19T04:47:39-0500

for simple interest:

"A=P(1+rt)"

"A=1350(1+0.045)=Php\ 1410.75"


it is added to the loan:

"A-P=1410.75-1350=Php\ 60.75"